OpenStudy (anonymous):

The sequence -2-11-26-47-74...-674 is neither arithmetic or geometric. How do I figure out the pattern to write it in summation notation?

OpenStudy (reemii):

first thing to do is looking at the difference between the terms. It is here: 11-2 = 9 26-11 = 15 47-26= 21 74-47=27 ... You notice that the differences grow by a +6 term.. After 74 it should be the number N such that N-74 = 27+6 = 33.

OpenStudy (reemii):

that means, the sequence is 2 = 2 2+9 = 11 11+9+6 = 26 26+9+6+6 = 47 47+9+6+6+6 = 74

OpenStudy (reemii):

write out a little more, it might give something.

OpenStudy (reemii):

Trick: look at 74. How many 9's in it? How many 6's in it? What about the next term? (107)

OpenStudy (reemii):

74 is the 5th term. There are 5-1 "9"'s in the 5th term. There are n-1 nines in the n-th term. There are 1+2+3 6's in the 5th term. There are 1+2+3+...+(n-2) sixes in the n-th term. And there's 2. So a_n = 2 + 9(n-1) + [1+2+...+(n-2)] * 6

OpenStudy (reemii):

And it seems to be correct.

OpenStudy (reemii):

by the way, you must use the formula for "1+2+...+k", with k=n-2 in order to obtain the real formula for \(a_n\).